If a is a root of 4x²+2x-1=0, then prove the other root is 2a2-2a.

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
a is the root of above equation say b is another root.
so 4a2+2a-1=0
a+b =-1/2
so put a=(-1/2-b)
4a2+2(-1/2-b)-1=0
4a2-1-2b-1=0
4a2-2b=2
2a2-b=1
2a2-1=b